# Words as Trigger Points in Social Media Discussions

### v1

@@ 3. Time
3.4 Features
We annotate several features to examine our initial hypotheses and study the impact of trigger words in online discussions. We simply use post counts for the first hypothesis (H1) related to user engagement. The second hypothesis (H2) requires a measure of animosity. To identify relevant features that can operationalise animosity, we build on various language models that follow the RoBERTa architecture (Liu et al. 2019) and are trained on social media corpora, specifically X (Twitter). Even though posts on Reddit tend to be longer than on X, the posts on both platforms share many similarities, e.g., the use of slang, emojis, unstructured text. Hence, we expect the trained models to perform well in either context (Priya et al. 2019).
User engagement (Hypothesis 1)
@@ 3.4 Features
Animosity (Hypothesis 2)
Hypothesis 2 suggests that trigger words cause more animosity in a debate. We operationalise this with controversial comments, the presence a negative sentiment, anger, and different forms of hate speech.
@@
Figure 2: Distribution of negatively charged comments for the treatment and control (space) groups of Vaccine as a trigger word.
@@
For sentiment analysis, we rely on the model twitter-roberta-base-sentiment-latest (Loureiro et al. 2022) which identifies a post as negative, neutral, or positive. This model has been fine-tuned for sentiment analysis using the SemEval 2017 Sentiment Analysis dataset (Rosenthal, Farra, and Nakov 2019). The sentiment in social media content has often been used to gain insights into the information related to controversial topics such as vaccination (Melton et al. 2021), politics (Guimaraes et al. 2019), and gender biases (Marjanovic, Stańczak, and Augenstein 2022). We expect trigger words to cause more negative sentiments.
@@
We rely on twitter-roberta-base-emotion-multilabel-latest (Camacho-Collados et al. 2022) to assign one or more emotions to each tweet. This model is trained with data from the Semeval 2018 Affect in Tweets task (Mohammad et al. 2018), covering 11 different emotions [^5] Detailed list of emotions can be found in Appendix A.1). For our analysis we focus on anger specifically, and we expect to observe more anger after a trigger word.
Figure 3: Distribution of comments in the treatment and control groups (space) for the trigger word Rwanda.
@@
Finally, we use the twitter-roberta-base-hate-multiclass model (Antypas and Camacho-Collados 2023) to detect hate speech. It is trained on a combination of 13 different hate speech Twitter datasets and can predict hate towards seven target groups [^6] Detailed list of categories can be found in Appendix A.2). The inclusion of hate speech detection as a feature is motivated by previous research revealing hateful trends in topics such as politics (Rieger et al. 2021), women’s rights (Farrell et al. 2019), and health issues (Walter et al. 2022). Considering the connotations of some of the triggers, we specifically focus on sexism for Feminism/Feminist, racism for Rwanda, while we consider all hateful comments jointly for NHS, Brexit and Vaccine.
4 Results
@@
4.1 User engagement (Hypothesis 1)
@@
When examining comment frequency across various control groups, the impact of specific trigger words becomes more apparent. For instance, Figure 3 illustrates the distribution of comments related to the trigger Rwanda in the space related experiment. It shows the overall number of messages across all threads. This trigger word leads to varying levels of engagement across different Reddit communities, with a notable increase in comments in identified communities where Rwanda is more controversial. Similar patterns are also observed in the semantic and temporal control groups for Rwanda (Appendix B, Figure 9).
@@
Table 5: Difference in Difference (DiD) scores along with the 95% confidence intervals (CI) for each setting tested. Bold scores indicate statistically significant results.
4.2 Animosity (Hypothesis 2)
@@
Figure 4: Average controversiality per thread for for the treatment and control (space) groups of Rwanda.
@@
We analyse the average controversiality of each thread and compare the proportion of controversial threads before and after the introduction of the trigger words. Our results (Table 5) indicate that, in general, there is an increase in controversial comments following each trigger word. Most terms display a statistically significant increase of DiD estimates when considering controversial messages in the treatment and control groups with the biggest difference observed on the semantic comparison for Rwanda, with a causal effect of 6.6% points.
Figure 5: Distribution of anger and non-anger comments for the treatment and control (space) groups of Vaccine trigger word.
@@
Figure 4 shows the percentages of controversial threads related to the trigger word Rwanda in both the space control and treatment groups, before and after its occurrence. The data indicates a noticeable increase in thread controversiality caused by the trigger word, reflected by a DiD estimate of 2.3% points. A similar trend is seen in both the semantic and temporal control settings, as documented in Appendix B, Figure 8).
@@
Figure 6: Distribution of hateful comments in the treatment and control groups (Space) of the trigger points Rwanda and Feminism.
@@
When considering the negatively charged comments of the treatment and control groups for each trigger word, Table 5 indicates a clear trend. We observe a significant increase in comments conveying a negative sentiment after the trigger word appears in 11 of the 12 settings tested, with the sole exception being the semantics comparison for NHS. The biggest effect in sentiment, with a DiD estimate of 5.5% points, appears for the Rwanda trigger word within both the space and semantics control groups.
Let us consider the vaccine trigger, Figure 2, as an example. Here, we again plot the overall number of messages we observe. Both treatment and space control groups exhibit a similar pattern of negative sentiment density as time approaches the trigger word, with an increase in negative sentiment just before and after the trigger word is mentioned. However, the treatment group shows more messages than the control group, suggesting that introducing the trigger words affects the expression of negative sentiment.
Similarly to the increase in negative sentiment, trigger words systematically cause more anger. We observe a positive causal effect in all conditions, with only 3 non-significant effects. Substantively, we measure a causal effect that ranges from 1.4% points more anger (Feminist) to 5.5% points (Vaccine semantic comparison). As another visual example, the overall number of posts expressing anger in the Vaccine treatment and space control groups increase significantly following the appearance of the trigger word, as shown in Figure 5. This is further confirmed by the DiD estimates, which indicate a significant difference in anger-related comments in 9 out of 12 cases.
@@
As a final indicator of animosity, we also consider hate speech. In our analysis, we use any kind of hate speech for Brexit, NHS, and Vaccine, and the more domain-specific racism-related hate speech for Rwanda and sexism-related hate speech for Feminism. As seen from Table 5, the impact of trigger words regarding the occurrence of hateful comments is not as prevalent as anger and negative sentiment. We observe significant effects in those areas where we can count on a domain-specific hate speech measure, i.e., in the cases of Feminism (space and semantic comparison) and Rwanda (space group).
@@
We exemplify our findings again in a visualisation, where we consider all messages irrespective of the thread they are posted in. We observe a distinct pattern, particularly in the context of Rwanda and Feminism where the volume of hateful content increases post-trigger. Figure 6 reveals the prevalence of racist and sexist comments within the treatment groups when compared to their control counterparts in regards to space.
5 Conclusion
@@
In this paper, we introduce the concept of trigger words, and study how they affect social media conversations. We reported empirical evidence based on a large-scale study on Reddit using five trigger words (Rwanda, Feminism, Brexit, NHS and Vaccine). In short, we find that trigger words cause more user engagement, but that this activity is marked by higher levels of animosity.
@@
Our findings confirm the first Hypothesis (H1), which suggests more user interaction after the appearance of a trigger word. We consistently observe more posts for all trigger words and across all control settings. Regarding our second hypothesis (H2), which concerns the impact of trigger words on the animosity of the ensuing discussions, we examined four different features that can indicate such effects. Controversiality, negative sentiment, and anger all show strong evidence that trigger words influence the discussion along those lines. We also investigated the causal effect of trigger words on hate speech. Where we could rely on a model that is capable of identifying domain-specific forms of hate speech—Sexism in the case of Feminism and Racism in the case of Rwanda—we find evidence that suggest trigger words play a causal role, here, too.
6 Limitations
@@
Despite potentially more accurate results, we refrained from using large language models (LLMs) for feature extraction mainly due to the vast size of our dataset (>>>100 million entries). Using LLMs would have significantly increased the computational and financial cost and the ecological impact of our experiments. Moreover, existing literature indicates that the zero-shot capabilities of LLMs do not consistently outperform smaller models for specific social media tasks like ours (Antypas et al. 2023).
@@
Selecting words as trigger points and the subreddits for analysing them offers room for improvement. The chosen subreddits may not fully represent the broader range of discussions these trigger words could provoke across the diverse Reddit community.
The goal of this paper is to verify trigger words as a concept. As such, we identify and manually select a few words as candidates according to our own expertise and social science literature. If we do not find evidence for the effects of trigger words in these easy cases, it would not make sense to study trigger words and their effects any further.
7 Ethics Statement
@@
All the data utilised in our research is sourced from publicly available information or collected using the official Reddit API. All the information is provided in an aggregated fashion, without reporting sensitive information from individual users.
The automatic methods used can be abused by those with the power to repress people with opposing opinions. Hence, we offer a comprehensive study but do not share our data publicly to avoid potential misuse or commercial use. Researchers interested in replicating our experiments can contact us directly.
@@ A.1 Emotion Categories
A.2 Hate Speech Categories.
@@ Appendix B Extended Results
Figure 7: Absolute count of posts before and after each trigger point.
@@
Figure 8: Average controversiality per thread in the Temporal and Semantic settings and their respective treatment groups, for the trigger point Rwanda.
@@
Figure 9: Rwanda Treatment and Control distributions of comments in the Semantic and Temporal settings.
@@
Figure 10: Distribution of negatively charged comments in the Treatment and Control (Space) groups.
Appendix C Computational Resources

### v2

@@ 3.3.2. Statistical analysis
Since the DiD estimator subtracts the difference before and after intervention in the control from the equivalent in the treatment group, no further control variables are necessary—if the parallel trends assumption holds (Roth et al., 2023). It requires that in the absence of treatment, the difference between the outcomes of the treated and control groups would follow the same trajectory over time. We investigate the plausibility of the assumption in section 4.3.
3.4. Features
We rely on several features to examine our initial hypotheses and study the impact of trigger words in online discussions. We simply use post counts for the first hypothesis (H1) related to user engagement. The second hypothesis (H2) requires a measure of animosity. To identify relevant features that can operationalise animosity, we build on various language models that follow the RoBERTa architecture (Liu et al., 2019) and are trained on social media corpora, specifically X (Twitter). Even though posts on Reddit tend to be longer than on X, the posts on both platforms share many similarities, e.g., the use of slang, emojis, or unstructured text. Hence, we expect the trained models to perform well in either context (Priya et al., 2019).
3.4.1. User engagement (Hypothesis 1)
@@
3.4.2. Animosity (Hypothesis 2)
Hypothesis 2 suggests that trigger words cause more animosity in a debate. We operationalise this with controversial comments, the presence of negative sentiment, anger, and different forms of hate speech.
@@ 3.4.2. Animosity (Hypothesis 2)
Figure 2. Distribution of negatively charged comments for the treatment and control (space) groups of Vaccine as a trigger word.
@@
For sentiment analysis, we rely on the model twitter-roberta-base-sentiment-latest (Loureiro et al., 2022) which identifies a post as negative, neutral, or positive. This model has been fine-tuned for sentiment analysis using the SemEval 2017 Sentiment Analysis dataset (Rosenthal et al., 2019). The sentiment in social media content has often been used to gain insights into the information related to controversial topics such as vaccination (Melton et al., 2021), politics (Guimaraes et al., 2019), and gender biases (Marjanovic et al., 2022). We expect trigger words to cause more negative sentiments.
@@
We rely on twitter-roberta-base-emotion-multilabel-latest (Camacho-Collados et al., 2022) to assign one or more emotions to each tweet. This model is trained with data from the Semeval 2018 Affect in Tweets task (Mohammad et al., 2018), covering 11 different emotions [^6] Detailed list of emotions can be found in Appendix A.1). For our analysis, we focus on anger specifically, and we expect to observe more anger after a trigger word.
Figure 3. Distribution of comments in the treatment and control groups (space) for the trigger word Rwanda.
@@
Finally, we use the twitter-roberta-base-hate-multiclass model (Antypas and Camacho-Collados, 2023) to detect hate speech. It is trained on a combination of 13 different hate speech Twitter datasets and can predict hate towards seven target groups [^7] Detailed list of categories can be found in Appendix A.2). The inclusion of hate speech detection as a feature is motivated by previous research revealing hateful trends in topics such as politics (Rieger et al., 2021), women’s rights (Farrell et al., 2019), and health issues (Walter et al., 2022). Considering the connotations of some of the triggers, we specifically focus on sexism for Feminism/Feminist, racism for Rwanda, while we consider all hateful comments jointly for NHS, Brexit and Vaccine.
4. Results
@@
4.1. User engagement (Hypothesis 1)
@@
When examining comment frequency across various control groups, the impact of specific trigger words becomes more apparent. For instance, Figure 3 illustrates the distribution of comments related to the trigger Rwanda in the space-related experiment. It shows the overall number of messages across all threads. This trigger word leads to varying levels of engagement across different Reddit communities, with a notable increase in comments in identified communities where Rwanda is more controversial. Similar patterns are also observed in the semantic and temporal control groups for Rwanda (Appendix B, Figure 9).
@@
Table 5. Difference in Difference (DiD) scores along with the 95% confidence intervals (CI) for each setting tested. Bold scores indicate statistically significant results.
4.2. Animosity (Hypothesis 2)
@@
Figure 4. Average controversiality per thread for the treatment and control (space) groups of Rwanda.
@@
We analyse the average controversiality of each thread and compare the proportion of controversial threads before and after the introduction of the trigger words. Our results (Table 5) indicate that, in general, there is an increase in controversial comments following each trigger word. Most terms display a statistically significant increase of DiD estimates when considering controversial messages in the treatment and control groups, with the biggest difference observed on the semantic comparison for Rwanda—a causal effect of 6.6% points.
@@
Figure 4 shows the percentages of controversial threads related to the trigger word Rwanda for the space experiment, before and after its occurrence. The data indicates a noticeable increase in thread controversiality caused by the trigger word, reflected by a DiD estimate of 2.3% points. A similar trend is seen in both the semantic and temporal control settings, as documented in Appendix B, Figure 8).
@@
Figure 5. Distribution of hateful comments in the treatment and control groups (space) of the trigger points Rwanda and Feminism.
@@
When considering the negatively charged comments of the treatment and control groups for each trigger word, Table 5 indicates a clear trend. We observe a significant increase in comments conveying a negative sentiment after the trigger word appears in 11 of the 12 settings tested, with the sole exception being the semantics comparison for NHS. The largest effect in sentiment, with a DiD estimate of 5.5% points, appears for the Rwanda trigger word in the space and semantics experiments.
Let us consider the vaccine trigger, Figure 2, as an example. Here, we again plot the overall number of messages we observe. Both treatment and space control groups exhibit a similar pattern of negative sentiment density as time approaches the trigger word, with an increase in negative sentiment just before and after the trigger word is mentioned. However, the treatment group shows more messages than the control group, suggesting that the trigger words affects the expression of negative sentiment.
Similarly to the increase in negative sentiment, trigger words systematically cause more anger. We observe a positive causal effect in all conditions, with only 3 non-significant effects. Substantively, we measure a causal effect that ranges from 1.4% points more anger (Feminist) to 5.5% points (Vaccine semantic comparison). As another visual example, the overall number of posts expressing anger in the Vaccine treatment and space control groups increase significantly following the appearance of the trigger word, as shown in Figure LABEL:fig:emotion_space_vaccine. This is further confirmed by the DiD estimates, which indicate a significant difference in anger-related comments in 9 out of 12 cases.
@@
As a final indicator of animosity, we also consider hate speech and abusive language, more specifically any kind of hate speech or abusive language for Brexit, NHS, and Vaccine, and the more domain-specific racism-related hate speech for Rwanda and sexism-related hate speech for Feminism. As seen from Table 5, the impact of trigger words regarding the occurrence of hateful or offensive comments is not as prevalent as anger and negative sentiment. We observe significant effects in those areas where we can count on a domain-specific hate speech measure, i.e., in the cases of Feminism (space and semantic comparison) and Rwanda (space group).
@@
We illustrate our findings again in a visualisation, where we consider all messages irrespective of the thread they are posted in. We observe a distinct pattern, particularly in the context of Rwanda and Feminism, where the volume of hateful content increases after the trigger. Figure 5 reveals the prevalence of racist and sexist comments within the treatment groups when compared to their control group in the space experiment.
4.3. Testing the Parallel Trends Assumption.
The DiD estimator relies on the assumption of parallel trends: without the treatment, the treated and control groups would follow the same trajectory over time. While there are no formal guarantees for the assumption to hold, there are means to argue in its favour. The first argument is a theoretical one. In neither of our experimental designs, we have reason to believe that the frequency of expressing animosity would not follow the same trend without the trigger word. Whether comparing treatment and control groups based on different Reddit fora (spaces), different words (semantics), or points in time (temporal), we suggest that without the respective trigger word, the relative difference between treatment and control group would remain the same.
We implement event studies for each experiment to further empirically support our argument. The central idea is to model the respective outcome variable as it unfolds over time, which allows us to assess the dynamics of treatment impact before and after the intervention. Plotting the treatment effects at different time points before the treatment, the event study helps visualise whether the pre-treatment trends in the treatment and control groups are indeed parallel. If the trends are not parallel before the intervention, the parallel trends assumption is violated, indicating potential bias in the DID estimator (Roth et al., 2023).
Figure 6. Testing the Parallel Trends Assumption with event studies. Time-wise mean of all individual event studies in purple.
For each of our overall 60 experiments, we divide the 30 minutes before and after the first mention of a trigger word into six distinct time intervals $\{[-30,-20[,[-20,-10[,[-10,0[,[0,10[,[10,20[,[20,30[\}${ [ - 30 , - 20 [ , [ - 20 , - 10 [ , [ - 10 , 0 [ , [ 0 , 10 [ , [ 10 , 20 [ , [ 20 , 30 [ } instead of just two. This allows us to follow, in particular, the pre-treatment trends at a higher granularity. We calculate each discussion thread’s share of (hostile) messages in the six intervals. Averaging over all these data points per period, we subtract the control group from the treatment group and calculate the difference of each observation from the first observation period.
Figure 6 charts the time-wise means over all 60 event studies. If the parallel trends assumption holds, the lines before the treatment at time = 0 should all remain on the horizontal line where the treatment effect = 0. Indeed, this is largely what we find. The mean treatment effect is close to 0 before the trigger word is mentioned for the first time. After the intervention, it is clearly above the horizontal 0-line. [^8] See Appendix for figures that include individual event studies for all experiments from H1 and H2.
5. Conclusion
@@
In this paper, we present the concept of trigger words and study how they affect social media conversations in UK politics. We report empirical evidence based on a large-scale study on Reddit using five trigger words (Rwanda, Feminism, Brexit, NHS and Vaccine). In short, our results show that trigger words cause more user engagement, but that this activity is marked by higher levels of animosity.
@@
Our findings confirm the first Hypothesis (H1), which suggests more user interaction after the appearance of a trigger word. We consistently observe more posts across all trigger words and control settings. Regarding our second hypothesis (H2), which concerns the impact of trigger words on the animosity of the ensuing discussions, we examine four different features that can indicate such effects. Trigger words cause more controversiality, more negative sentiment, and more anger. We also investigate the causal effect of trigger words on hate speech. Where we can, we rely on a model that is capable of identifying domain-specific forms of hate —sexism in the case of Feminism and racism in the case of Rwanda—. Our findings similarly confirm that trigger words play a causal role in this aspect as well.
6. Limitations
@@
We refrain from using large language models (LLMs) for feature extraction mainly due to the vast size of our dataset (>>>100 million entries). Using LLMs would have significantly increased the computational and financial cost and the ecological impact of our experiments. Moreover, existing literature indicates that the zero-shot capabilities of LLMs do not consistently outperform smaller models for specific social media tasks like ours (Antypas et al., 2023).
@@
Selecting words as trigger points and their corresponding subreddits offers room for improvement. The chosen subreddits may not fully represent the broader range of discussions these trigger words could provoke across the diverse Reddit community. Similarly, we only investigate a very select set of potential trigger words in the context of the UK.
However, our study had different goal: We consider our work to be a pilot study that intends to investigate whether the trigger points exist in online settings at all. Based on the empirical evidence we find, we can be confident to have found relevant empirical evidence that warrants further investigation.
7. Ethics Statement
@@
All data utilised in our research is sourced from publicly available information or collected using the official Reddit API. All information is provided in an aggregated fashion without reporting sensitive information from individual users.
@@ A.1. Emotion Categories
A.2. Hate Speech Categories.
@@ Appendix B Extended Results
Figure 7. Absolute count of posts before and after each trigger point.
@@
Figure 8. Average controversiality per thread in the Temporal and Semantic settings and their respective treatment groups, for the trigger point Rwanda.
@@
Figure 9. Rwanda Treatment and Control distributions of comments in the Semantic and Temporal settings.
@@
Figure 10. Distribution of negatively charged comments in the Treatment and Control (Space) groups.
Appendix C Extended Test of Parallel Trends Assumption
(a) Experiments from H1.
(b) Experiments from H2.
Figure 11. Testing the Parallel Trends Assumption with event studies. Individual studies in grey, their time-wise mean in purple.
Appendix D Computational Resources